OpenStudy (mony01):

A box with an open top is to be constructed from a square piece of cardboard, 3 ft wide, by cutting out a square from each of the four corners and bending up the sides. Find the largest volume that such a box can have.

OpenStudy (anonymous):



OpenStudy (anonymous):

So this is a figure of what your box is going to look like.

OpenStudy (anonymous):

Volume = (3-x)(3-x)(x)

OpenStudy (anonymous):

Domain (0,3)

OpenStudy (anonymous):

Graph

OpenStudy (anonymous):

My bad volume = (3-2x)(3-2x)(x)

OpenStudy (anonymous):

MAximum volume is 27

OpenStudy (mony01):

What would be the expression for the volume V in terms of x and y.

OpenStudy (anonymous):

y = (3-2x)(3-2x)(x)

OpenStudy (mony01):

it says its wrong

OpenStudy (anonymous):

What does it give you

OpenStudy (mony01):

Write an expression for the volume V in terms of x and y. Use the given information to write an equation that relates the variables x and y. Use it to write the volume as a function of x. Finish solving the problem by finding the largest volume that such a box can have.

OpenStudy (mony01):

for each one i have to write an answer

OpenStudy (anonymous):

I'm not sure then. Sorry